OpenStudy (anonymous):

Which number supports the given conjecture? If the sum of the digits of a number is divisible by 9, then the number is also divisible by 9. A. 19 B. 54 C. 87 D. 109

OpenStudy (anonymous):

So this is saying that you need to add up the digits and see if it is divisible by nine. So for example if the number they gave you was 27 you would add 2+7 = 9 and divide by 9. Since you get a whole number (not a fraction or decimal) that would be correct. Do that with the numbers you were given.

OpenStudy (anonymous):

so the answer is 19

OpenStudy (anonymous):

No, it's no. 9+1=10. Which is not divisible by 9.

OpenStudy (anonymous):

so is it 54

OpenStudy (anonymous):

Are you just guessing or are you actually trying this out?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

oh so its 109

OpenStudy (anonymous):

Why do you think it's 109? What happened when you followed the steps to prove it?

OpenStudy (anonymous):

i think its 109 because 9+1=10 and 10 is not divisible by 9 so i guessed 109 because 10 and 9

OpenStudy (anonymous):

It is not 109. 1+0+9 = 10. 10/9 is 1.11111. What you need is a final result that is a whole number. Again, refer to my example of 27. 2+7=9. 9/9=1 a whole number.

OpenStudy (anonymous):

87

OpenStudy (anonymous):

If you think it's 87, prove why.

OpenStudy (anonymous):

cause i ask u if it was 109 and u said no and i ask u if it was 54 and u didnt say nothing and u said it is not 19

OpenStudy (anonymous):

I've told you how to solve this. I'm not going to give you the answer if you're not going to try to solve it properly.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

You didn't, you just guessed. You've narrowed it down to 54 or 87 because I didn't explicitly say either of those were wrong. What happens when you apply the steps I told you to follow to those numbers?

OpenStudy (the_fizicx99):

Just take the numbers apart, add them and see if it's divisible by 9.

OpenStudy (the_fizicx99):

By "divisible" they mean it has no remainder, so you can't have a decimal. Rather a whole number.